Status Report

NASA Space Station On-Orbit Status 8 Jul 2003

By SpaceRef Editor
July 8, 2003
Filed under , , ,

All ISS systems continue to function nominally, except as noted
previously or below.  Day 75 in space for the crew.  

Using the MO-21 "Ecosfera" air sampler and incubation
equipment, broken out and set up yesterday, CDR Malenchenko spent about
40 minutes collecting air samples for atmospheric microbial
[MO-21 determines microbial contamination of the ISS
atmosphere, specifically the total bacterial and fungal microflora counts
and microflora composition according to morphologic criteria of
microorganism colonies.]

Malenchenko then worked on restoring the Russian laptop 1 to
functionality by first creating a boot-up floppy disk (for Norton Ghost
5.1D) on Russian laptop 2.  Later, he activated laptop 1, installed
a software patch and "cloned" a new HDD (hard disk drive) on
the machine from the Norton Ghost 5.1D disk obtained from laptop 2.

FE/SO Ed Lu meanwhile spent time in the U.S. segment (USOS) with the
regular periodic inspection of smoke detectors in the Lab module, Airlock
(A/L) and Node "Unity", as well as of the A/L’s microbial

Yuri Malenchenko completed the scheduled installation of new satellite
navigation equipment (ASN), delivered on Progress 11P, in the Service
Module (SM). 
[The ASN was connected to the onboard computer
system’s laptop 3, configured for the subsequent test.  After he had
mated BITS2-12 onboard measurement telemetry system connectors to the
ASN, a raw data transfer test to laptop 3 was to be conducted from TsUP
by remote control.]

The CDR attended to the routine maintenance of SOZh life support systems,
while Ed Lu prepared the daily IMS (inventory management system)
"delta" file for updating the IMS database and completed the
regular daily status checkup of the autonomous Lab payloads..

Ed also attended to the weekly transfer of accumulated data files with
workout data from the TVIS and RED exercise equipment to the MEC via
PCMCIA cards (personal computer memory card international association)
for subsequent downlink to Earth. 

At 5:45am EDT, both crewmembers participated in an interactive televised
PAO/educational event sponsored by NASDA, with Japanese middle school
students gathered at the Tokyo FM Radio station banquet hall (Ku- &
S-band).  Former NASDA Astronaut Mamoru Mohri attended.

Both crewmembers completed their daily physical exercise program (2.5
hrs.) on TVIS treadmill, CEVIS cycle ergometer, and RED anaerobic

The Science Officer changed lenses of the EarthKAM (EK) camera at the Lab
nadir window from 50mm to 180mm.  EK team reported an excellent
startup yesterday, and the on-board system is working well.  

[Attending today’s and tomorrow’s National Geographic
Society-sponsored sessions are 25 students, ranging from

grade to college sophomores. 
Commented one of their teachers:  "I can’t think of another
program that interests and excites students to the point that they are
willing to come in during vacation (some home from college), to work
during a mission."  EarthKAM (Earth Knowledge Acquired by
Middle school students) employs an ESC 460C electronic still camera
mounted at the Lab’s down-facing window, normally equipped with a 50mm
lens (f1.4).  Powered by 16Vdc from a 120 Vdc adapter, EarthKAM
takes pictures by remote operation from the ground, without crew
interaction.  It is available for Middle School-ers, who submit
image requests and conduct geographic research.  The requests are
uplinked in a camera control file to an SSC (station support computer)
laptop which then activates the camera at specified times and receives
the digital images from the camera’s storage card on its hard drive, for
subsequent downlink via OpsLAN.]

The Kodak 760 ECS (electronic still camera) used by Ed Lu to
photograph stunning aurora images has shown some noise, due to
"pixelation" or "hot pixels" not properly removed or
blended into the image by the automatic camera procedure.  The most
probable cause is excessive radiation as ISS passed through the
aurora.  All noise will be removed from the photos by ground

Today’s CEO targets, no longer limited in the current LVLH
attitude and including the targets of the Lewis & Clark 200-year
memorial locations, were
Baghdad, Iraq

(nadir pass),

Phoenix, AZ

(NIGHT TARGET: nadir pass),


(nadir pass),

Pomeroy, Washington

& CLARK SITE:  Nadir pass–looking on the south side of the
Snake River [within the second bend of the river upstream of its
confluence with the Columbia River]),

Lemhi Pass,

(LEWIS & CLARK SITE:  Nadir pass–shooting
mountain range crests [dark forest green, as opposed to lighter, browner
river valleys] and the crew should have gotten this now-unused pass over
the continental divide.  The explorers crossed here into Idaho on
August 12, 1805.  Lemhi Pass was abandoned for the westward
migrations since other lower passes exist),


(nadir pass), and

Los Angeles, California

(nadir and a touch left).

CEO images can be viewed at the websites


See also the website "Space Station Challenge"

U.S. and Russian Segment Status

(as of 1:30pm

Environmental Control and Life Support (ECLSS) and Thermal Control

Elektron O2 generator is powered On (16 amp mode).  Vozdukh
CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is
Off.  TCCS (trace contaminant control subsystem) is operating. 
MCA (major constituents analyzer) is operating.  BMP Harmful
Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode
(unit is off).  RS air conditioner SKV-1 is Off; SKV-2 is Off.

SM Working Compartment:
  Pressure (mmHg) — 742; temperature
(deg C) — 27.5; ppO2 (mmHg) — data invalid
; ppCO2 (mmHg) —
data invalid

SM Transfer Compartment:
  Pressure (mmHg) — 751; temperature
(deg C) — 20.4.

FGB Cabin:
  Pressure (mmHg) — 752; temperature (deg C) —

  Pressure (mmHg) — 741.34; temperature (deg C) — 23.0
(shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

U.S. Lab:
  Pressure (mmHg) — 743.40; temperature (deg C) —
n/a; ppO2 (mmHg) — 169.1; ppCO2 (mmHg) — 3.8.

Joint Airlock (Equip. Lock):
  Pressure (mmHg) — 743.50;
temperature (deg C) — n/a; shell heater temp (deg C) — 22.3, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.

  Shell heater temp (deg C) — 23.9

  Shell heater temp (deg C) — 20.2.

         (n/a = data not

Propulsion System (PS):

Total propellant load available:
3803 kg (8384 lb)

of 6/26 
[SM(774) + FGB(2447) + Progress M(182) +Progress
M-1(400)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

Both P6 channels fully operational.  BGA (beta gimbal
assembly) 2B and 4B both in directed position (Blind/triple-angle mode,
non solar-tracking, drag reduction)..
SM batteries:  Battery #4 is disconnected (failed 6/16); batteries
#2, 6 and 8 are in "Cycle" mode; all other batteries (4) are in
"Partial Charge" mode.  Batteries #1, #2 and #3 showing
degraded capacity. 

FGB batteries:  Battery #3 is offline; battery #1 is in
"Cycle" mode; all other batteries (4) are in "Partial
Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in
GNC-1 MDM is prime; GNC-2 is Backup.

INT-1 is operating; INT-2 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).

LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is Off; PL-1 MDM is Operational.

APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.

SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).

State vector source — U.S. SIGI-1 (GPS)
Attitude source — U.S. SIGI-1 (GPS)

Angular rate source — RGA-1

Flight Attitude:

LVLH -YVV (local vertical/local horizontal =
"earth-fixed": z-axis in local vertical, -y-axis in velocity
vector [yaw: -10 deg, pitch: -9.3 deg, roll: 0 deg]), with CMG/TA
(thruster assist) Momentum Management.

Solar Beta angle:  24.6 deg (magnitude increasing).

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.

All other Russian communications & tracking systems are
S-band is operating nominally (on string 2).

Ku-band is operating nominally.
Audio subsystem is operating nominally [IAC-1 (internal audio controller
#1) being analyzed after self-test error.  IAC-2 is prime.]

Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating


SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power
on both strings.
MBS: KA power on both strings. 

MT: latched and mated at WS4. 
POA: KA power on both strings.

RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is

ISS Orbit
  (as of this morning, 8:06am EDT [=

Mean altitude — 386.9 km

Apogee — 392.3 km
Perigee — 381.4 km

Period — 92.29 min.
Inclination (to Equator) —  51.63 deg

Eccentricity — 0.0008085
Orbits per 24-hr. day — 15.60

Solar Beta Angle — 24.6 deg (magnitude increasing)
Mean altitude loss in last 24 hours — 75 m

Revolutions since FGB/Zarya launch (Nov. ’98)  — 26433
For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.